The tangent line to the graph of y = h(x) at the point (-1, 7) passes through the point (9, 9). Find h(-1) and h'(-1).
Answer by Boreal(15235)

h(-1)=7
point slope formula y-y1=m(x-x1), m slope, (x1,y1) point
slope is 2/10=1/5
y-9=(1/5)(x-9)=(1/5)x-(9/5)
y=(1/5)x+(36/5) the equation of the tangent line.
The slope of the tangent line is the derivative of h(x). That is the slope of the line which is (1/5). The previous -1/5 was in error.